# EVOS: Efficient Implicit Neural Training via EVOlutionary Selector

### v1

@@ Abstract
We propose EVOlutionary Selector (EVOS), an efficient training paradigm for accelerating Implicit Neural Representation (INR). Unlike conventional INR training that feeds all samples through the neural network in each iteration, our approach restricts training to strategically selected points, reducing computational overhead by eliminating redundant forward passes. Specifically, we treat each sample as an individual in an evolutionary process, where only those fittest ones survive and merit inclusion in training, adaptively evolving with the neural network dynamics. While this is conceptually similar to Evolutionary Algorithms, their distinct objectives (selection for acceleration vs. iterative solution optimization) require a fundamental redefinition of evolutionary mechanisms for our context. In response, we design sparse fitness evaluation, frequency-guided crossover, and augmented unbiased mutation to comprise EVOS. These components respectively guide sample selection with reduced computational cost, enhance performance through frequency-domain balance, and mitigate selection bias from cached evaluation. Extensive experiments demonstrate that our method achieves approximately 48%-66% reduction in training time while ensuring superior convergence without additional cost, establishing state-of-the-art acceleration among recent sampling-based strategies.
@@ 2.2 Acceleration for INR Training
Numerous approaches have been proposed to address the computational intensity of INR training, which can be categorized into four main streams: (1) Partition-based methods decompose the input signal into smaller components, employing multiple compact MLPs for local representation. These include approaches based on regular blocks [43], adaptive blocks [30], Voronoi diagrams [42], Laplacian pyramids [45], and segmentation maps [26]. (2) Prior-driven methods leverage knowledge distilled from large-scale datasets to expedite convergence. This category encompasses meta-learning approaches for parameter initialization [54, 10] and hypernetwork-based weight generation [49, 51]. (3) Explicit caching methods trade spatial complexity for temporal efficiency by maintaining encoded features in explicit data structures, such as hash maps [33, 56], point cloud [57], low-rank tensors [6], and tree structures [58, 27]. (4) Alternative approaches explore diverse strategies to enhance training efficiency, including modulator networks [31], reparameterized training [48], data transformation [47], and batch normalization [4].
@@ 3.3 Frequency-Guided Crossover
Since the $L_{lap}(\cdot)$ operator requires a full volume of attributes for its spatial convolution filter, the sparsified coordinates $\mathbf{z}_{t}$ cannot be directly applied. We address this by incorporating cached evaluation results $\hat{F}_{\theta}(\mathbf{x}\setminus\mathbf{z}_{t})$ from the last key iteration. $\lambda_{l}$ and $\lambda_{h}$ denote the respective scaling factors.
@@ 4.1 Implement Details
Evaluation Mechanisms. We evaluated reconstruction quality using PSNR, SSIM [55], and LPIPS [61] metrics. The efficiency improvements were assessed through two mechanisms: (1) comparing time required to achieve target reconstruction quality, and (2) comparing performance and computational cost under fixed iterations. The first approach directly demonstrates efficiency gains across different strategies, while the second reveals the per-iteration performance difference between full data utilization and partial data sampling. All reported results were averaged across the dataset, with each experiment repeated five times to ensure statistical reliability.
@@ 5 Conclusion
In this paper, we propose EVOlutionary Selector (EVOS) to accelerate INR training. Inspired by natural selection, EVOS selects only the fittest coordinates for network optimization, reducing redundant computations while enhancing training efficiency. Our method comprises three key modules: sparse fitness evaluation, frequency-guided crossover, and augmented unbiased mutation, which enable efficient coordinate selection, enhance optimization performance, and mitigate selection bias. Extensive experiments demonstrate EVOS’s superior performance, surpassing existing sampling-based acceleration methods. Notably, we discover that strategic sparsification of training samples not only reduces computational cost but also improves per-iteration performance, suggesting that sample weighting techniques can enhance optimization in signal fitting tasks beyond their traditional role in model generalization.

### v3

@@ Abstract
We propose EVOlutionary Selector (EVOS), an efficient training paradigm for accelerating Implicit Neural Representation (INR). Unlike conventional INR training that feeds all samples through the neural network in each iteration, our approach restricts training to strategically selected points, reducing computational overhead by eliminating redundant forward passes. Specifically, we treat each sample as an individual in an evolutionary process, where only those fittest ones survive and merit inclusion in training, adaptively evolving with the neural network dynamics. While this is conceptually similar to Evolutionary Algorithms, their distinct objectives (selection for acceleration vs. iterative solution optimization) require a fundamental redefinition of evolutionary mechanisms for our context. In response, we design sparse fitness evaluation, frequency-guided crossover, and augmented unbiased mutation to comprise EVOS. These components respectively guide sample selection with reduced computational cost, enhance performance through frequency-domain balance, and mitigate selection bias from cached evaluation. Extensive experiments demonstrate that our method achieves approximately 48%-66% reduction in training time while ensuring superior convergence without additional cost, establishing state-of-the-art acceleration among recent sampling-based strategies. Our code is available at this link.
@@ 2.2 Acceleration for INR Training
Numerous approaches have been proposed to address the computational intensity of INR training, which can be categorized into four main streams: (1) Partition-based methods decompose the input signal into smaller components, employing multiple compact MLPs for local representation. These include approaches based on regular blocks [43], adaptive blocks [30], Voronoi diagrams [42], Laplacian pyramids [45], and segmentation maps [26]. (2) Prior-driven methods leverage knowledge distilled from large-scale datasets to expedite convergence. This category encompasses meta-learning approaches for parameter initialization [55, 10] and hyper-network-based weight generation [50, 52]. (3) Explicit caching methods trade spatial complexity for temporal efficiency by maintaining encoded features in explicit data structures, such as hash maps [33, 57], point cloud [58], low-rank tensors [6], and tree structures [59, 27]. (4) Alternative approaches explore diverse strategies to enhance training efficiency, including modulator networks [31], reparameterized training [49], data transformation [47, 64], batch normalization [4], and gradient adjustment [48].
@@ 3.3 Frequency-Guided Crossover
Since the $L_{lap}(\cdot)$ operator requires a full volume of attributes for its spatial convolution filter, the sparsified coordinates $\mathbf{z}_{t}$ cannot be directly applied. We address this by incorporating cached evaluation results $\hat{F}_{\theta}(\mathbf{x}\setminus\mathbf{z}_{t})$ from the last key iteration. $\lambda_{l}$ and $\lambda_{h}$ denote the scaling factors.
@@ 4.1 Implement Details
Evaluation Mechanisms. We evaluated reconstruction quality using PSNR, SSIM [56], and LPIPS [62] metrics. The efficiency improvements were assessed through two mechanisms: (1) comparing time required to achieve target reconstruction quality, and (2) comparing performance and computational cost under fixed iterations. The first directly reveals efficiency gains across different strategies, while the second compares the per-iteration performance between full data utilization and partial data sampling.
@@ 5 Conclusion
In this paper, we propose EVOlutionary Selector (EVOS) to accelerate INR training. Extensive experiments demonstrate EVOS’s superior performance, surpassing existing sampling-based acceleration methods. Notably, we discover that strategic sparsification of training samples not only reduces computational cost but also improves per-iteration performance, suggesting that sample weighting techniques can enhance optimization in signal fitting tasks beyond their traditional role in model generalization.
Acknowledgments
This work is supported in part by National Key Research and Development Project of China (Grant No. 2023YFF0905502), National Natural Science Foundation of China (Grant No. 92467204 and 62472249), and Shenzhen Science and Technology Program (Grant No. JCYJ20220818101014030 and KJZD20240903102300001). We thank anonymous reviewers for their valuable advice.